# Guided Policy Optimization under Partial Observability

### v1

@@ 3.2 GPO-Penalty
This section introduces a straightforward implementation of the GPO framework using KL-divergence as a penalty for the guider, which we refer to as GPO-penalty. Specifically, in step 2 of GPO, we use PPO as the underlying trust-region algorithm. The corresponding objective for the guider’s policy is as follows [^1] We omit subscripts for expectations in the remainder of the paper, as all samples are drawn from the distribution induced by the behavioral policy $\beta=\mu_{\text{old}}$.:
$\mathcal{L}_{1}(\mu)=\mathbb{E}\bigg{[}\min\bigg{(}\rho^{\mu}A^{\beta}(s,a),\rho_{clip}^{\mu}A^{\beta}(s,a)\bigg{)}\bigg{]},$ (2)
where $\rho^{\mu}=\mu(a|s)/\beta(a|s)$, $\rho_{clip}^{\mu}=clip(\rho^{\mu},1-\epsilon,1+\epsilon)$ and $\beta$ denotes the behavioral policy. The advantage $A^{\beta}(s,a)$ is estimated using the Generalized Advantage Estimation (GAE) (Schulman2015HighDimensionalCC, ) with the value function $V(s)$ trained via discounted reward-to-go.
@@
A key insight in GPO is that exact backtracking of the guider’s policy is unnecessary—it’s sufficient to keep the guider within a "possibly good" region relative to the learner. The learner may fail to follow the guider either because the guider is inimitable or just because the guider learns faster, the latter being common due to inexact gradient updates. In such cases, aggressive backtracking can be harmful. Keeping the guider slightly ahead also allows it to collect better trajectories, as discussed in Section 4.4. To maintain this balance, we introduce a coefficient $\alpha$ that modulates the guider’s objective as
$\mathcal{L}(\mu)=\mathcal{L}_{1}(\mu)-\alpha\mathcal{L}_{3}(\mu),$ (5)
@@
This scheme is analogous to the KL-penalty adjustment in PPO-penalty (DBLP:journals/corr/SchulmanWDRK17, ), where the penalty coefficient adjusts based on the relationship between the KL divergence and a predefined threshold.
Another key aspect is compensating for the learner’s policy improvement, as we replace strict backtracking with a KL constraint. While it is possible to set a very small $d_{\text{targ}}$, this would inefficiently inflate $\alpha$, hindering the guider’s training. Notably, Proposition 1 implies that applying GPO with PPO is effectively equivalent to applying PPO directly to the learner. Consequently, we can concurrently train the learner’s policy using PPO during the GPO iterations. As a result, we introduce an additional objective for the learner’s policy:
$\mathcal{L}_{4}(\pi)=\mathbb{E}\bigg{[}\min\bigg{(}\rho^{\pi}A^{\beta}(s,a),\rho_{clip}^{\pi}A^{\beta}(s,a)\bigg{)}\bigg{]},$ (7)
@@
For policy $\pi$, $\mu$, $\beta$ and all states $s$, suppose $D_{TV}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$, then we have
@@
$\mathcal{L}(\pi)=\alpha\mathcal{L}_{4}(\pi)-\mathcal{L}_{2}(\pi),$ (8)
where the coefficient $\alpha$ from (6) is applied to the RL term. This mechanism compensates when the learner struggles to follow the guider. If the learner is able to fully track the guider, $\alpha$ approaches zero, allowing the guider to directly lead the learner to the optimal policy without requiring an additional RL objective. When the learner cannot keep pace, the RL objective aids in the learner’s training.
@@ 3.3 GPO-Clip
In this section, we introduce a slightly modified implementation of the GPO framework, which we refer to as GPO-clip. The key principle is that an effective guider should remain at the boundary of the learner’s “possibly good" region: if the guider is too far ahead, the learner struggles to follow; if too close, the guider’s ability to provide effective supervision and better trajectory diminishes. To achieve this balance, the guider should halt updates when it moves too far ahead and avoid backtracking when it is already sufficiently close.
We propose two key modifications to the GPO-penalty algorithm introduced in the previous subsection. First, inspired by PPO-clip, we replace the clip function $\rho_{clip}^{\mu}$ in (2) with the following double-clip function:
@@
It is important to note that, unlike PPO where PPO-clip can completely replace the KL-penalty term, this is not the case in GPO. In PPO, the ratio $\rho^{\pi}(s,a)$ starts at 1 at the beginning of each epoch, ensuring that the clipped ratio keeps $\pi$ near the behavioral policy. In GPO, however, the gap between $\pi(a|s)$ and $\mu(a|o)$ may accumulate over multiple updates if the learner fails to keep up with the guider. The double-clip function (9) alone is insufficient to bring $\pi(a|o)$ back into the $\delta$ region once it has strayed too far. To address this, we introduce a mask on the backtracking loss, defined as:
$m(s,a)=\mathbb{I}\big{(}\frac{\pi(a|s)}{\mu(a|o)}\notin(1-\delta,1+\delta)\big{)},$ (10)
@@
$L_{\text{GPO-penalty}}(\theta)=$ $\mathbb{E}\Big{[}\min\Big{(}\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big{)}-\alpha\text{D}_{\text{KL}}\big{(}\mu_{\theta}(\cdot|o_{g})||\pi_{\hat{\theta}}(\cdot|o_{l})\big{)}$ (11)
$\alpha\min\Big{(}\rho^{\pi_{\theta}}A^{\beta}(o_{l},a),\rho_{clip}^{\pi_{\theta}}A^{\beta}(o_{l},a)\Big{)}-\text{D}_{\text{KL}}\big{(}\mu_{\hat{\theta}}(\cdot|o_{g})||\pi_{\theta}(\cdot|o_{l})\big{)}\Big{]},$
$L_{\text{GPO-clip}}(\theta)=$ $\mathbb{E}\Big{[}\min\Big{(}\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta},\pi_{\hat{\theta}}}A^{\beta}(o_{g},a)\Big{)}-m(s,a)\text{D}_{\text{KL}}\big{(}\mu_{\theta}(\cdot|o_{g})||\pi_{\hat{\theta}}(\cdot|o_{l})\big{)}$ (12)
$\alpha\min\Big{(}\rho^{\pi_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\pi_{\theta}}A^{\beta}(o_{g},a)\Big{)}-\text{D}_{\text{KL}}\big{(}\mu_{\hat{\theta}}(\cdot|o_{g})||\pi_{\theta}(\cdot|o_{l})\big{)}\Big{]},$
@@ 4 Experiments
In this section, we evaluate the empirical performance of GPO across various domains. Section 4.1 presents didactic tasks to verify GPO’s properties, such as optimality. Section 4.2 evaluates GPO on partially observable and noisy continuous control tasks in the Brax (brax2021github, ) environment, comparing it against several baselines. Section 4.3 evaluates GPO’s performance on memory-based tasks from POPGym (morad2023popgym, ), and Section 4.4 provides ablation studies and further discussion.
@@
For baselines, we consider two types of approaches for utilizing teacher supervision. The first type involves training both the teacher and student simultaneously. A summary of their main characteristics is provided in Table 1. Among these, GPO-naive refers to GPO-penalty without the RL auxiliary loss. PPO-asym directly trains the learner using PPO, with the learner’s value function receiving $o_{g}$ as input. PPO+BC trains the teacher with PPO, while the learner is trained via direct BC from the teacher. ADVISOR-co is a modification of ADVISOR (10.5555/3540261.3541724, ), and A2D is based on the work by (Warrington2020RobustAL, ). The second type involves training the teacher first, followed by the application of TSL methods. These include DAgger (ross2011reduction, ), PPO+BC-t, ADVISOR, ELF (Walsman2023ImpossiblyGE, ), and ELF-asym, where ELF is a policy distillation method that utilizes reward shaping to provide supervision to the student, and ELF-asym is a variant that uses an asymmetric value function. Further details about these algorithms can be found in Appendix E.1.
@@ 4.1 Didactic Tasks
We begin by evaluating our algorithm on two didactic problems introduced in Section 2.2. As shown in Fig. 2, direct cloning of the guider’s policy converges to a suboptimal solution, as expected. In contrast, all variants of GPO achieve optimal performance on these tasks. Although applying RL directly to the learner easily leads to optimal solutions, it is important to note that GPO-naive achieves optimality purely through supervised learning. This result verifies the optimality guarantee of the GPO framework described in Proposition 1, suggesting that a guider constrained within the learner’s “possibly good" region can provide effective supervision, even with asymmetric information. Moreover, comparing GPO-naive to GPO-penalty and GPO-clip reveals that the introduction of direct RL training for the learner accelerates learning.
@@ 4.2 Continuous Control Tasks in Brax
In this subsection, we present the results of our algorithms and baselines on several continuous control tasks in the Brax domain. To transform these tasks into a POMDP setting, we remov the velocity information of all joints, and add varying levels of noise to the observations. The guider has access to full, noiseless information, while the learner operates with partial and noisy inputs. For more details, please refer to Appendix E.
Figure 3: Comparison of GPO and baselines on the Brax domain, where $\sigma$ represents the scale of Gaussian noise added to the observations. The performance on each task is normalized to $[0,1]$ using the performance of the corresponding pre-trained teacher as a reference. Algorithms highlighted in red are supervised by their corresponding pre-trained teacher.
The results are shown in Fig. 3, where the performance hierarchy is generally: GPO-clip $>$ GPO-penalty $>$ PPO-asym $>$ GPO-naive $>$ other baselines. It is important to note that, even without factoring in the cost of training the teacher (which is nearly the same as training the GPO algorithm from scratch), methods that rely on a pre-trained privileged teacher perform well only in the Halfcheetah and Swimmer tasks. Furthermore, the performance of these methods declines rapidly as the noise scale increases. This occurs because, when the pre-trained teacher becomes too skilled for the student, it provides little to no useful supervision, and may even have a negative impact on learning.
For co-training approaches, we have the following key observations: First, the superior performance of GPO-clip and GPO-penalty compared to the base algorithm PPO shows that this framework can effectively utilize additional information during training to facilitate the learner training. Second, comparing GPO-naive to GPO-penalty and GPO-clip, we see that introducing RL training for the learner improves performance. Third, the comparison between PPO+BC and GPO-naive highlights the necessity of backtracking. If the guider is not constrained to the learner, the guider’s supervision may negatively influence the performance. Last, other baselines such as ADVISOR failed to utilize the privileged teacher such that it degenerates into pure PPO.
In summary, our method consistently outperforms the baselines, demonstrating its effectiveness in solving noisy and partially observable continuous control tasks. Other baselines, such as L2T-RL (wu2024learnteachimprovesample, ) and TGRL (shenfeld2023tgrl, ), are not presented here as they are not directly comparable; however, additional experiments are provided in Appendix E.4.
@@ 4.3 Memory-based Tasks in POPGym
Fig. 4 shows the results on 15 POPGym tasks, where we compare GPO-penalty and GPO-clip to PPO-asym and PPO. The general conclusion mirrors the results from the previous subsection, where GPO-clip typically outperforms GPO-penalty, followed by PPO-asym and PPO. Key insights include: First, the superior performance of GPO-penalty indicates that the ability of the guider to explore further without diverging too much from the learner proves valuable in these memory-based tasks. Second, while PPO-asym outperforms PPO, its performance improvement is less pronounced here than in the Brax domain, suggesting that asymmetric value function may not be very helpful for memory tasks. Third, although neither GPO-penalty nor GPO-clip exhibits superior performance in tasks like BattleshipMedium and CountRecallHard, this is due to we use the same parameter across all tasks, and performance could be improved as we show in the next section.
@@ 4.4 Ablations and Discussions
Why do GPO outperform other baselines? We attribute the superior performance to two factors: effective RL training of the learner, and effective supervision from the guider. The benefit of RL training is shown in Fig. 6(left), where GPO-ablation (GPO-penalty without supervision, as described in Table 1) outperforms PPO-asym on the Humanoid task. Although both use similar objectives, GPO-ablation uses data collected by the guider, indicating that a better behavior policy improves learning efficiency. The effectiveness of the supervision comes from the guider being constrained to the “possibly good" region while still learning rapidly. In Fig. 6(right), with the learner trained purely by supervision (GPO-clip with RL disabled), GPO-clip outperforms GPO-ablation, PPO+BC, and PPO-asym. This shows that in memory-intensive tasks, supervision is more beneficial than RL. Since PPO+BC performs poorly in noisy tasks in Section 4.2 but comparably to PPO-asym here, we can also infer that supervision plays a particularly important role in these tasks. Moreover, GPO-clip’s strong performance over PPO+BC—despite both using pure supervision—highlights the importance of constraining the guider to a policy the learner can follow.

### v2

@@ 3.2 GPO-Penalty
This section introduces a straightforward implementation of the GPO framework using KL-divergence as a penalty for the guider, which we refer to as GPO-penalty. Specifically, in step 2 of GPO, we use PPO as the underlying trust-region algorithm. The corresponding loss for the guider’s policy is as follows [^1] We omit subscripts for expectations in the remainder of the paper, as all samples are drawn from the distribution induced by the behavioral policy $\beta=\mu_{\text{old}}$.:
$\mathcal{L}_{1}(\mu)=-\mathbb{E}\bigg[\min\bigg(\rho^{\mu}A^{\beta}(s,a),\rho_{clip}^{\mu}A^{\beta}(s,a)\bigg)\bigg],$ (2)
where $\rho^{\mu}=\mu(a|s)/\beta(a|s)$, $\rho_{clip}^{\mu}=clip(\rho^{\mu},1-\epsilon,1+\epsilon)$ and $\beta$ denotes the behavioral policy. The advantage $A^{\beta}(s,a)$ is estimated using the Generalized Advantage Estimation (GAE) (Schulman et al., 2015b) with the value function $V(s)$ trained via discounted reward-to-go.
@@
A key insight in GPO is that exact backtracking of the guider’s policy is unnecessary—it’s sufficient to keep the guider within an imitable region relative to the learner. The learner may fail to follow the guider either because the guider is inimitable or because the guider learns faster, the latter being common due to inexact gradient updates. In such cases, aggressive backtracking can be harmful. Keeping the guider slightly ahead also allows it to collect better trajectories, as discussed in Section 4.4. To maintain this balance, we introduce a coefficient $\alpha$ that modulates the guider’s loss as
$\mathcal{L}(\mu)=\mathcal{L}_{1}(\mu)+\alpha\mathcal{L}_{3}(\mu),$ (5)
@@
This scheme is analogous to the KL-penalty adjustment in PPO-penalty (Schulman et al., 2017), where the penalty coefficient adjusts based on the relationship between the KL divergence and a predefined threshold.
Another aspect is compensating for the learner’s policy improvement, as we replace strict backtracking with a KL constraint. While it is possible to set a very small $d_{\text{targ}}$, this would inefficiently inflate $\alpha$, hindering the guider’s training. Notably, Proposition 1 implies that applying GPO with PPO is effectively equivalent to applying PPO directly to the learner. Consequently, we can concurrently train the learner’s policy using PPO during the GPO iterations. As a result, we introduce an additional objective for the learner’s policy:
$\mathcal{L}_{4}(\pi)=-\mathbb{E}\bigg[\min\bigg(\rho^{\pi}A^{\beta}(s,a),\rho_{clip}^{\pi}A^{\beta}(s,a)\bigg)\bigg],$ (7)
@@
For policy $\pi$, $\mu$, $\beta$ and all states $s$, suppose $\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$, then we have
@@
$\mathcal{L}(\pi)=\alpha\mathcal{L}_{4}(\pi)+\mathcal{L}_{2}(\pi),$ (8)
where the coefficient $\alpha$ from equation 6 is applied to the RL term. This mechanism compensates when the learner struggles to follow the guider. If the learner is able to fully track the guider, $\alpha$ approaches zero, allowing the guider to directly lead the learner to the optimal policy without requiring an additional RL objective. When the learner cannot keep pace, the RL objective aids in the learner’s training.
@@ 3.3 GPO-Clip
In this section, we introduce a slightly modified implementation of the GPO framework, which we refer to as GPO-clip. The key principle is that an effective guider should remain at the boundary of the learner’s imitable region: if the guider is too far ahead, the learner struggles to follow; if too close, the guider’s ability to provide effective supervision and better trajectory diminishes. To achieve this balance, the guider should halt updates when it moves too far ahead and avoid backtracking when it is already sufficiently close.
We propose two modifications to the GPO-penalty algorithm introduced in the previous subsection. First, inspired by PPO-clip, we replace the clip function $\rho_{clip}^{\mu}$ in equation 2 with the following double-clip function:
@@
It is important to note that, unlike PPO where PPO-clip can completely replace the KL-penalty term, this is not the case in GPO. In PPO, the ratio $\rho^{\pi}(s,a)$ starts at 1 at the beginning of each epoch, ensuring that the clipped ratio keeps $\pi$ near the behavioral policy. In GPO, however, the gap between $\pi(a|s)$ and $\mu(a|o)$ may accumulate over multiple updates if the learner fails to keep up with the guider. The double-clip function equation 9 alone is insufficient to bring $\pi(a|o)$ back into the $\delta$ region once it has strayed too far. To address this, we introduce a mask on the backtracking loss, defined as:
$m(s,a)=\mathbb{I}\big(\frac{\mu(a|s)}{\pi(a|o)}\notin(1-\delta,1+\delta)\big),$ (10)
@@
$L_{\text{GPO-penalty}}(\theta)=\mathbb{E}\Big[$ $-\min\Big(\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big)+\alpha\text{D}_{\text{KL}}\big(\mu_{\theta}(\cdot|o_{g}),\pi_{\hat{\theta}}(\cdot|o_{l})\big)$ (11)
$-\alpha\min\Big(\rho^{\pi_{\theta}}A^{\beta}(o_{l},a),\rho_{clip}^{\pi_{\theta}}A^{\beta}(o_{l},a)\Big)+\text{D}_{\text{KL}}\big(\mu_{\hat{\theta}}(\cdot|o_{g}),\pi_{\theta}(\cdot|o_{l})\big)\Big],$
$L_{\text{GPO-clip}}(\theta)=\mathbb{E}\Big[$ $-\min\Big(\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta},\pi_{\hat{\theta}}}A^{\beta}(o_{g},a)\Big)+m(s,a)\text{D}_{\text{KL}}\big(\mu_{\theta}(\cdot|o_{g}),\pi_{\hat{\theta}}(\cdot|o_{l})\big)$ (12)
$-\alpha\min\Big(\rho^{\pi_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\pi_{\theta}}A^{\beta}(o_{g},a)\Big)+\text{D}_{\text{KL}}\big(\mu_{\hat{\theta}}(\cdot|o_{g}),\pi_{\theta}(\cdot|o_{l})\big)\Big],$
@@ 4 Experiments
In this section, we evaluate the empirical performance of GPO across various domains. For baselines, we consider two types of approaches for utilizing teacher supervision. The first type involves training both the teacher and student simultaneously. A summary of their main characteristics is provided in Table 3. Among these, GPO-naive refers to GPO-penalty without the RL auxiliary loss. PPO-asym directly trains the learner using PPO,
@@
with the learner’s value function receiving $o_{g}$ as input. PPO+BC trains the teacher with PPO, while the learner is trained via direct BC from the teacher. ADVISOR-co is a modification of ADVISOR (Weihs et al., 2024), and A2D is based on the work by (Warrington et al., 2020). The second type involves training the teacher first, followed by the application of TSL methods. These include DAgger (Ross et al., 2011), PPO+BC-t, ADVISOR, ELF (Walsman et al., 2023), and ELF-asym, where ELF is a policy distillation method that utilizes reward shaping to provide supervision to the student, and ELF-asym is a variant that uses an asymmetric value function. Further details about these algorithms can be found in Appendix E.1.
@@ 4.1 Didactic Tasks
We begin by evaluating our algorithm on two didactic problems introduced in Section 2.2. As shown in Fig. 2, direct cloning of the guider’s policy converges to a suboptimal solution, as expected. In contrast, all variants of GPO achieve optimal performance on these tasks. Although applying RL directly to the learner easily leads to optimal solutions, it is important to note that GPO-naive achieves optimality purely through supervised learning. This result verifies the optimality guarantee of the GPO framework described in Proposition 1, suggesting that a guider constrained within the learner’s imitable region can provide effective supervision, even with asymmetric information. Moreover, comparing GPO-naive to GPO-penalty and GPO-clip reveals that the introduction of direct RL training for the learner accelerates learning.
@@ 4.2 Continuous Control Tasks in Brax
In this subsection, we present the results of our algorithms and baselines on several continuous control tasks in the Brax domain. To transform these tasks into a POMDP setting, we remove the velocity information of all joints, and add varying levels of noise to the observations. The guider has access to full, noiseless information, while the learner operates with partial and noisy inputs. For more details, please refer to Appendix E.
Figure 3: Comparison of GPO and baselines on the Brax domain, where $\sigma$ represents the scale of Gaussian noise added to the observations. The performance on each task is normalized to $[0,1]$ using the performance of the corresponding pre-trained teacher as a reference. Algorithms highlighted in red rely on a separately pre-trained privileged teacher.
The results are shown in Fig. 3, where the performance hierarchy is generally: GPO-clip >> GPO-penalty >> PPO-asym >> GPO-naive >> other baselines. It is important to note that, even without factoring in the cost of training the teacher (which is comparable to training GPO from scratch), methods that rely on a pre-trained privileged teacher perform well only in the Halfcheetah and Swimmer tasks. Furthermore, the performance of these methods declines rapidly as the noise scale increases. This occurs because, when the pre-trained teacher becomes too skilled for the student, it provides little to no useful supervision, and may even have a negative impact on learning.
For co-training approaches, we have the following observations: First, the superior performance of GPO-clip and GPO-penalty compared to the base algorithm PPO shows that this framework can effectively utilize additional information during training to facilitate the learner training. Second, comparing GPO-naive to GPO-penalty and GPO-clip, we see that introducing RL training for the learner improves performance. Third, the comparison between PPO+BC and GPO-naive highlights the necessity of backtracking. If the guider is not constrained to the learner, the guider’s supervision may negatively influence the performance. Last, other baselines such as ADVISOR failed to utilize the privileged teacher such that it degenerates into pure PPO.
In summary, our method consistently outperforms the baselines, demonstrating its effectiveness in solving noisy and partially observable continuous control tasks. Additional experiments including L2T-RL (Wu et al., 2024), TGRL (Shenfeld et al., 2023b) and RMA (Kumar et al., 2021) are provided in Appendix E.4.
@@ 4.3 Memory-based Tasks in POPGym
Fig. 4 shows the results on 15 POPGym tasks, where we compare GPO-penalty and GPO-clip to PPO-asym and PPO. The general conclusion mirrors the results from the previous subsection, where GPO-clip typically outperforms GPO-penalty, followed by PPO-asym and PPO. Key insights include: First, the superior performance of GPO-penalty indicates that the ability of the guider to explore further without diverging too much from the learner proves valuable in these memory-based tasks. Second, while PPO-asym outperforms PPO, its performance improvement is less pronounced here than in the Brax domain, suggesting that asymmetric value function may not be very helpful for memory tasks. Third, although neither GPO-penalty nor GPO-clip exhibits superior performance in tasks like BattleshipMedium and CountRecallHard, this is due to the fact that we use the same parameter across all tasks, and performance could be improved as we show in the next section.
@@ 4.4 Ablations and Discussions
Why do GPO outperform other baselines? We attribute the superior performance to two factors: effective RL training of the learner, and effective supervision from the guider. The benefit of RL training is shown in Fig. 6(left), where GPO-ablation (GPO-penalty without supervision, as described in Table 3) outperforms PPO-asym on the Humanoid task. Although both use similar objectives, GPO-ablation uses data collected by the guider, indicating that a better behavior policy improves learning efficiency. The effectiveness of the supervision comes from the guider being constrained to the imitable region while still learning rapidly. In Fig. 6(right), with the learner trained purely by supervision (GPO-clip with RL disabled), GPO-clip outperforms GPO-ablation, PPO+BC, and PPO-asym. This shows that in memory-intensive tasks, supervision is more beneficial than RL. Since PPO+BC performs poorly in noisy tasks in Section 4.2 but comparably to PPO-asym here, we can also infer that supervision plays a particularly important role in these tasks. Moreover, GPO-clip’s strong performance over PPO+BC—despite both using pure supervision—highlights the importance of constraining the guider to a policy the learner can follow.